OpenStudy (pig_cat223):

plz help... solve this system of equations. seperate the x- and y-values with a comma. 20x+11y=37 4x+3y=13

OpenStudy (theprofessor35):

First, use substitution.

OpenStudy (theprofessor35):

20x+11y=37 4x+3y=13

OpenStudy (theprofessor35):

20x+11y=37 11y = 37 - 20x

OpenStudy (theprofessor35):

11y = 37 - 20x y = 37/11 - 20/11x

OpenStudy (theprofessor35):

plug that into the other equation

OpenStudy (theprofessor35):

4x+3y=13 y = 37/11 - 20/11x 4x + 3(37/11 - 20/11x) = 13